Displaying reports 401-420 of 1081.Go to page Start 17 18 19 20 21 22 23 24 25 End
Reports until 15:24, Monday 14 October 2024
GQuEST General
Daniel Grass - posted 15:24, Monday 14 October 2024 (11903)
Parameters Figured out to make Breadboard Pattern in 304(L) Stainless Steel

I spent some time figuring out how to get the HAAS VF-2 CNC Machine to make high quality holes into 304(L) Stainless Steel. Specifically, I wanted to make 2 0.625" deep holes with a 1/4-20 tap 0.5" deep and connected with a 1/8" groove to avoid virtual leaks.

For the pilot hole, the standard diameter for a 1/4-20 hole is 0.201". This seemed to break the tap, so I used a 0.209" diameter drill that Paul Stovall, the MCE machinist, recommended. This means the threads are a bit weaker, but they will be plenty strong for our purposes. Specifically, I used a #4 gauge carbide drill bit with 2 flutes and a TiAlN coating.

To drill into the material, I used a spindle speed of 2300 rpm and a plunge feed rate of 7 in/min. This is a bit slower than what the manufacturer recomended (11 in/min plunge rate) but in line with what FS Wizard calculated. 11 in/min broke the drill bit. Importantly, after each hole I spun the drill bit backwards at 2000 rpm. This removes the chips that can get stuck on the drill and break it. I also used coolant through the spindle to help with chips.

After drilling, I used a 3/8" diameter 90° chamfer tool to chamfer the holes to make threading easier and with fewer burrs. I used a spindle speed of 1000 rpm and a plunge feed rate of 2.3 in/min. To be safe with chips, I spun the chamfer tool backwards at 800 rpm. I used flood coolant.

I then used this 1/4-20 steel bottom tap with 3 flutes and a black oxide coating. A true bottom tap has a bit less of a chamfer; this tap is better for starting threads. This CNC machine can't use multiple taps. I used a spindle speed of 200 rpm. I did a chip breaking cycle ever 5 mm. I used coolant through the spindle.

Lastly, to add a groove, I used a 1/8" diameter carbide end mill with 4 flutes and a TiAlN coating. I used a spindle speed of 5136 rpm and a cutting feed rate of 4.76 in/min. I used flood coolant. To help with the longevity of the end mill, I plunged over the existing hole. This groove is only 0.015" deep to help with tool life.

I have added plenty of pauses after each drill to ensure the chips are gone. If the chips remain, I can stop the CNC machine.

Next, I will write some CAM software for the tapped flange and run it.

GQuEST General
Daniel Grass - posted 15:42, Friday 11 October 2024 (11900)
Shelf Placed in Mobile Clean Rooms

[Torrey, Sander, Alex, Daniel]

We placed the overhead shelf over the tables within the mobile clean rooms. The idea is to go between the different subsystems of GQuEST and go to the electronics rack.

We moved the shelf using the two furniture dollies.

See attached photo.

Images attached to this report
GQuEST General
Daniel Grass - posted 15:39, Friday 11 October 2024 (11899)
Shelf Placed in Mobile Clean Rooms

[Torrey, Sander, Alex, Daniel]

We placed the overhead shelf over the tables within the mobile clean rooms. The idea is to go between the different subsystems of GQuEST and go to the electronics rack.

See attached photo.

Atoms General
Briana Chen - posted 07:25, Wednesday 09 October 2024 (11766)
Are there more photons or atoms?

In an old post I was curious about whether it is possible there are too many photons bombarding atoms in the vapor cell, causing saturation. However, this doesn't seem to be too reasonable of an argument anymore because based on the attached rough calculations, there tends to be more Rb atoms by a few orders of magnitude. So, a better calculation probably deals with the absorption coefficient/lifetime of the transition.

Also, I have just now realized how the vapor cell actually works: there is some solid Rb and gaseous Rb (2-5 mg for Rb cells according to Thorlabs) at room temperature and when heated, more of the Rb evaporates so there is then a higher number density. So in these calculations it seems reasonable to use the vapor pressure since in this case Rb vapor is exerting pressure on the condensed phases.

Non-image files attached to this report
Atoms General
Daniel Grass - posted 17:18, Tuesday 08 October 2024 (11897)
PICAS (2D MOT) Supports Cleaned

I cleaned the 2D MOT Supports so that they can be near vacuum and good optics.

Vacuum Systems General
Daniel Grass - posted 17:17, Tuesday 08 October 2024 (11896)
Small Titanium Hex Keys cleaned to Class B UHV Standard

I cleaned a 5/64", 1/16", and 0.050" titanium hex keys for Class B UHV use. 

Electronics General
Torrey Cullen - posted 13:04, Tuesday 08 October 2024 - last comment - 21:56, Monday 14 October 2024(11894)
Electronics moving out of b102

[Jeff, Torrey]

We have started moving electronics out of B102. We started with the three Moku's not in use by the RBQ experiment. Two have been mounted in the rack as seen in the photo. The 3rd doesn't have a mount yet, Jeff is emailing liquid instruments about this.

Images attached to this report
Comments related to this report
Torrey Cullen - 13:09, Tuesday 08 October 2024 (11895)

I've disassembled and moved all main electronics associated with the filter cavities to B111. This includes the SHG temperature controller, 1811 15V power sources, RF amplifiers for AOMs (two free standing ones and the two in the chasis Daniel made), HV amplifier for piezo control, a spare function generator, and all 7 photo detector power sources. A permanent spot for this has not yet been found.

Jeffrey Wack - 16:07, Friday 11 October 2024 (11901)

Moku is sending us two sets of rack mounting hardware (for free)

Torrey Cullen - 14:21, Monday 14 October 2024 (11902)

[Jeff, Daniel, Torrey]

We have wrapped the optical sleds in Mylar and are ready to move. With Sander and Alex out, and Ian still injured, we need to wait til Daniel Jeff and I are all free at the same time to make the final move. Tentatively, this will be Tuesday. I may ask JC to join as well. Hopefully moving optics soon!

Jeffrey Wack - 21:56, Monday 14 October 2024 (11904)

All 4 Moku Pros now have rack mounting fins installed

GQuEST General (GQuEST, Layouts)
Sander Vermeulen - posted 11:58, Tuesday 08 October 2024 (11893)
Moved vacuum chambers onto optics tables in

We have moved one vertex vacuum chamber ('beamsplitter chamber'/'service vessel') onto the table into the final position for the 'first', i.e. the south-east-most IFO (see designed layout here: Link). In addition, we have moved two of the cube vacuum vessels (formerly the end mirror chambers of the Holometer) into the desired position to comprise the vacuum enclosure for the 'first' (i.e. the south) laser filter cavity.

The move was done with an engine hoist, where we suspended the chambers close to over their desired position, and performed the remaining rotational and horizontal positioning by torqueing and pulling on the suspended chamber from the sides, after which the chamber was slowly lowered. 

The rotational alignment of the vertex chamber was verified using a laser level, where we first aligned the laser beam along the flange that will define the future transmitted IFO arm by aligning the laser across two bolts along the diameter of the top flange by eye (see attached photo). We then checked the distance of the beam from the wall in B111D to verify alignment of the 'arm' with the room. We found the arm flange to be rotated from a line parallel to the wall by 1.4 degrees, which is probably within the error of our measurement, and is certainly within the standard deviation of the process of rotating the chamber when placing it down using the engine hoist. Importantly, this alignment will allow the arm to clear the the door frame in the intended position.

The east vacuum cube for the LFC was placed a couple of inches further away from the west cube than it will be for the final configuration, to allow the other vacuum components to be placed between them.

Images attached to this report
GQuEST End Mirror Mount (GQuEST)
Daniel Grass - posted 19:38, Monday 07 October 2024 (11891)
Knight Optical Silicon Optics Surface Microroughness Measurements

[Lee, Daniel]

We worked with Albert Lazzarini to use an instrument to measure the microroughness of the surface of the Knight Optical silicon optics. We were able to record images of the surface with ~3 mm side length. We took measurements of optics with a cylindrical barrel and octagonal barrel in the center and close to the spoke. The data as screenshots and .csv is on the nextcloud, but it is down right now. It seems the optics are flat with an rms between 0.1 nm and 1 nm, which should be small enough for our purposes. A full analysis is needed. Note that the x and y axes (in the screenshots at least) is in units of 10 μm. This was determined by knowing how much we moved the stage and by tracking a piece of schmutz. We confirmed this by moving to the side of the optics. See attached photos. See below for instructions from Albert.

 

Open 4Sight Focus

Click Video for live view

Red: gain too high, so lower exposure

Also change gain

 

Get fringes with (tip/tilt and) focus

 

Select Autofocus

Select Coarse Autofocus

 

Repeat a few times

 

Play with gain and exposure so there’s no railing

Better to have exposures lower

 

Hit measure

 

Can also do burst for more (ended up doing this but took a few minutes)

 

Can save as csv

 

When done:

Close program

Log out

Turn off switch

Turn off computer if files are transferred

Images attached to this report
How To General
Briana Chen - posted 20:58, Sunday 06 October 2024 - last comment - 18:17, Sunday 20 October 2024(11820)
Get plant transfer function from laser locking

Following this tutorial: Laser Locking With Closed-Loop Transfer Function Measurement (liquidinstruments.com), here's a summary of useful points in measuring the transfer function of the plant, open loop system, etc, which can be useful in the design choices for the laser PID controller (like determining how good the gain parameters are). This was done with just the slow controller enabled. It would be a little more complicated if both slow and fast controllers are enabled because you would have to divide out the fast controller TF (from the Moku configuration, Out A of the LLB would not be the pure error signal). 

Note: there is an error in Figure 1 of the linked tutorial. Noise should be injected AFTER the plant for the subsequent math to be correct. 

Basic idea: To test how good your plant can deal with disturbances, you inject some disturbance and measure the system response. A way to measure these responses is shown here: sys_tf.jpg (we combined the sensing S into the plant P, but the same math applies to the tutorial). Essentially, the math suggests that if we inject disturbances and measure their corresponding output, we can get the open and closed loop transfer function of the system, which informs us about the stability of our controller.

We need to split up the feedback loop into the following components to do this: schematic.jpg. We rigged up the Moku in this configuration to measure all these parameters: Mokuconfig.jpg, using the Laser Lock Box (LLB), Lock In Amplifier (LIA), PID, and Frequency Response Analyzer (FRA). You move the demodulation process into the Lock-In Amplifier (LIA) and perform the injection using a PID as an adder. Then, you use a Frequency Response Analyzer to get the bode plots for the different transfer functions (open loop, plant, etc). In the FRA, you apply a frequency scan to measure the transfer functions. 

We locked the laser first using the following controller: slowcontroller.png. We measure the open loop transfer function (corrected for the phase going beyond 360 degrees), where we find unity gain (0 dB) occurs at 1.85 Hz with a phase margin of 65 deg: OpenLoopGainTransferFunction.pdf. This means that the system is reasonably stable (a large phase shift is needed to reach instability so the system is robust) and can control below 1.85Hz (can't suppress noise effectively above this).

This open loop TF is given by CP. We can divide CP by the controller transfer function (C) that we applied (dividing the gain and the phase) to get the plant transfer function, which is shown as a Bode plot here: PlantTransferFunction.pdf. A few notes:

Knowing the plant TF, you can use these functions to find potential resonant peaks or bandwidth limits in the system. This is useful because you want to make sure your controller won't start to amplify those peaks and drive the system out of stable lock. The plant transfer function bode plot has a generally flat magnitude, which is expected because it should not be amplifying any particular frequency. The slight bump at ~4 Hz may be some slight resonance in the system, it shows up in the final noise spectrum as well interestingly enough (NoiseSpectra.pdf), although this is with the slow and fast controller. Ideally, you want the range of your open loop phase margin to be between -90 and 90 degrees for as large of a controller bandwidth as possible. We would be able to model the system fully if we could sum the noise spectra (inside closed loop) and plant transfer function. 

We can also check the goodness of our controller by looking at the disturbance rejection, which is the response put on a disturbance going through the system. You would want the bode plot of the disturbance rejection to be at low magnitude (e.g. negative dB) to suppress the disturbance. To ensure that the perturbations are suppressed enough and you can hold the lock with less noise, the magnitude of the disturbance rejection should not increase beyond 0 dB after the unity gain frequency because then, noise will be amplified in the system (i.e. (1/(1+CP))*N = disturbance rejection TF * Noise --> increased noise in the output). I don't really understand the last sentence of this article where the low frequency gain of the open loop TF matching the disturbance rejection magnitude suggests there is adequate noise suppression: I would think that as long as the magnitude of 1/(1+CP) is small, the magnitude of CP should not matter. 

Images attached to this report
Non-image files attached to this report
Comments related to this report
Briana Chen - 18:17, Sunday 20 October 2024 (11916)

Plant transfer function pdf didn't work, here's the actual thing.

Non-image files attached to this comment
Atoms General
Daniel Grass - posted 21:07, Friday 04 October 2024 (11889)
PICAS (2D MOT) Supports Machined

I designed and machined some supports for the PICAS 2D MOT. These will support the MOT and allow it to be attached with the bellows to the chamber holding the 3D MOT to trap and cool the Rb atoms. See attached photo and part files. The part is not meant to clamp the 2.75" flange but go around it.

Images attached to this report
Non-image files attached to this report
Lab Infrastructure General
Alex Ramirez - posted 13:23, Wednesday 02 October 2024 (11888)
File path for Office Signs Illustrator File

Here is where to find the Illustrator files for the McCuller Lab signage in Bridge Basement.

\Nextcloud\Lab\signs_and_nametags\Room_Names_Template.ai

There is also a Word document for printing out more name tags if you need another for your desk. 
 

Please contact me, Alex, before saving or modifying the adobe illustrator files or save a new one as your own. 

For additional signs to be made, please get in touch with me on mattermost

Lasers Laser Filter Cavity (GQuEST)
Daniel Grass - posted 18:24, Tuesday 01 October 2024 (11887)
Custom Based and Internally Tapped 10 Inch Flange Attached to Laser Filter Cavity Cubes

[Alex, Ian, Jeff, Daniel]

I one by one took out the 12 point hex screws that secured the custom base of the Laser Filter Cavity (LFC) Cube and inserted and moderately tightened nickel coated hex screws, with a smaller screw head heights, and washers. I then did this process for the 12 point hex screws that directly secured the Tapped 10" Flange and added washers. I then tightened the screws in a star pattern until I could no longer see the copper gasket. See attached photos.

To put the cube on the base, we lifted from the same 2 lifting points in this post. The cube naturally tilts such that the base goes down furthest when lowered again.

We lifted the 2nd LFC Cube in the same way as the first and repeated the same process. This time, I used torque wrenches to tighten the screws, incrementing by 4-5 Nm every time. The ultimate tightness needed to no longer see the copper gasket was 33 Nm. I recommend going by more than 4-5 Nm becuase it adds more opportunity to strip a screw. I nearly stripped one, but was able to tighten it to 30 Nm. We then put the cube on its base.

 

Images attached to this report
Lab Infrastructure General
Daniel Grass - posted 12:23, Monday 30 September 2024 - last comment - 20:32, Monday 07 October 2024(11884)
B111A Oxygen Sensor

[Ian, Daniel, Sander]
 

Ian and I plugged in the East Bridge B111A O2 monitor, and it read 0.0 and then FLT then I unplugged it after a second when it alarmed.

 

Comments related to this report
Daniel Grass - 20:32, Monday 07 October 2024 (11892)

This seems to be fixed now.

Equipment Loans
Daniel Grass - posted 14:48, Monday 23 September 2024 - last comment - 12:56, Friday 11 October 2024(11875)
Silver Plated 12 Point Bolts Loaned to Paco

I loaned a bag of Silver Plated 12 Point Bolts (1/4-28 x 7/8") with 1/4" washers, 25/pk to Paco. Duniway Part Number SBX-28-087. Paco said he would return the bag on Thursday, 9/26.

Comments related to this report
Daniel Grass - 12:56, Friday 11 October 2024 (11898)

A new bag of this equipment has been traded to us.

Atoms General
Briana Chen - posted 16:20, Wednesday 28 August 2024 - last comment - 17:17, Sunday 06 October 2024(11840)
Pump updates as of 8/22

[Briana, Ian, Torrey]

Ian suggested that one reason why the tank circuit didn't work was because the capacitance is actually 12 not 14 picofarads based on the spec sheet that came in the EOM box (linked in a previous log post). This changes the inductance from 139 nH to 119 nH, which may be within the tolerance of the inductor we used but might still amplify the wrong frequency.

Thanks to Ian and Torrey for realigning the pump before I destroyed it. We are trying to mode-match the pump and probe beams through the vapor cell (concentrate more pump power to sharpen the dip), so using JamMt and beam profiling, two lenses (150 mm and 100 mm) were added to the pump path to reduce the beam size down from ~1600 microns to ~1300 microns. One issue is that once the light passed through the EOM, the beam profile began to look more clipped, so the x and y beam diameters were not the same and changed drastically. I think this is because the beam passing through the EOM is not constant in the z-direction, so more fine-tuning is required that I was not able to finish in time. By tuning the position of the EOM, the beam size can change significantly, so it is almost certainly the EOM alignment causing this issue. Currently, from the beginning to end of the vapor cell, the probe beam has a (x,y) beam profile of (1349, 1455) and (1405, 1545) microns respectively. For the pump, this was (1368, 1361) and (1383, 1386) microns at the beginning and end of the vapor cell respectively. 

Spent some time varying the slow controller and there exists a peak around 40 Hz introduced by the locked laser (potentially due to the controller but unsure otherwise): example_peak_pt2.pngstrangepeak.png show some different controller parameters. If you reduce the unity gain frequency and the gain, this peak gets damped down, which makes sense because the gain will be lower at that frequency. This peak could be a result of us driving the system at this frequency that is in resonance with something in the system, kind of like when you have a spring and you drive it at the right frequency to increase the amplitude but once you are at higher/lower frequencies to this peak you have a lower amplitude.

We achieved fast actuation by changing the current controller. We previously did not implement this because we thought power fluctuations might become too significant. We also failed to do so initially because we needed to amplify the signal to the current controller by +14 dB in the Moku settings. On 8/21, we took data with the fast controller but the error signals looked extremely bad: horrendous_dip.pngplateau_errorsignal.png. Although the calibrated error signals for the 8/22 data also looks ugly, this one is worse because the error signal strangely plateaus. After plugging in a terminator to the BNC cable and unplugging, this plateau no longer occurred. This points to the potential fluctuations in the current controller due to the new connection from the MOD OUT port with the BNC cable. Still, the data from 8/21 shows that the slow controller is doing something. By locking and then unlocking the slow controller, we get the noise spectrum on the right side of this image: slowvsfast.png. The following are the settings used to take this data: settings.png. The controllers used were shown here: slowcontroller.pngfastcontroller.png.

The following is the error signal used for the 8/22 data. When the signal is not calibrated, it looks pretty smooth: uncalibratederrorsignal.png. Upon calibrating, the overall error signal looks bad: uglyerrorsignals.png. This is not because of the sample acquisition rate. I am pretty sure this is because the current value may be fluctuating now that we have it connected to the BNC cable/Moku. What we can do to test this is to plug in the current output to a Moku input, convert that voltage reading to current, and perform the interpolation for current similar to the one we do with temperature using the laser specs. This calibrates using a potentially varying current value along with the varying temperature value. Doing this would require three data loggers though, which is not super efficient.

On 8/22, we retook data after fixing the error signal plateau-ing. We obtain the following noise spectra: Probe(withoutpump).pdfProbe(withpump).pdf, which matches what we expect with a lower noise in the locked laser. The controllers used to lock are: slowcontroller_822.pngfastcontroller_822.png. After calibrating with the error signal (which gave us slopes of ~3.4 V/THz and ~21.6 V/THz for the probe and probe with pump respectively, showing the improved frequency discrimination), we get the following noise spectra: NoiseSpectra.pdf. This seems good but we should keep the following in mind.

Some other to-dos:

Images attached to this report
Non-image files attached to this report
Comments related to this report
Briana Chen - 17:17, Sunday 06 October 2024 (11883)

Fixed issue with terrible calibration: the temperature scan signal was weirdly noisy (see CorrectedTempscan.jpg), so I did a linear fit to smooth it over and used these values to perform calibration. This led to updated calibration values for the slope of the linear region: Probe: Slope (V/nm):  -1.659, Slope (V/THz):  3.415 | Probe with pump: Slope (V/nm):  -13.004, Slope (V/THz):  26.2367

Attached are the fixed calibration graphs: err_signal_with_pump.pdferr_signal_without_pump.pdf. The updated noise spectra with these slopes is found here: NoiseSpectra.pdf

Images attached to this comment
Non-image files attached to this comment
Displaying reports 401-420 of 1081.Go to page Start 17 18 19 20 21 22 23 24 25 End